OpenStudy (anonymous):

Problem: You are helping design a water slide for children ranging in size from 35kg to 75kg. You have been tasked with creating the mats for the kids to sit on whilst they travel down the water slide. The slide's angle is 37 degrees with respect to the horizontal. For safety reasons, you can't let the children speed up any faster than dv/dt = 3.0 m/s2. What is the minimum coefficient of friction needed between the mat and the slide? (Hint: What does F_net need to be?) Work I've done: F_net = m*dv/dt (Approx.), F_net = 3.0m/s2 * 75kg = 225N.

OpenStudy (anonymous):

you are right about F_net; and for coefficient you need to solve \[F_{net}=mg\sin(\theta)-\mu mg\cos(\theta)\]